# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@ 5 Examples
Consider the Example 4.4 again. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ be the product measure, which is a $\sigma$-ergodic measure on $[0,1]^{\mathbb{Z}}$. Then the ergodicity of $\mu$ enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in[0,1]^{\mathbb{Z}}.$
Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ24, Theorem 1.1], we have
@@
We extend the Example 5.1 to more general full shift.
@@
Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
For $n\geq 0$, let $\pi_{n}$ be the projection from $K^{\mathbb{Z}}$ to its first $n$ coordinates.
It was proved in [S22, Theorem 15]
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K),$
where $\overline{\rm dim}_{B}(K)$ is the upper box dimension of $K$ under the metric $d$.
@@
Let $K$ be a compact metric space with a metric $d$. Then there exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
@@
It is known that the system $(K^{\mathbb{Z}},\sigma)$ has specification property. Let $S_{\epsilon}$ be a maximal $\epsilon$-separated subset of $K$ with respect to metric $d$. For $n\geq 1$, we put
${\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
and consider the set $\hat{\mathcal{S}}_{n}\subset K^{\mathbb{Z}}$ consisting of one preimage of each element in ${\mathcal{S}_{n}}$ under $\pi_{n}$. Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$. Define
$\mu_{n}:=\frac{1}{\#\hat{\mathcal{S}}_{n}}\sum_{x\in\hat{\mathcal{S}}_{n}}\mathcal{E}_{n}(x).$
We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(K^{\mathbb{Z}})$. Then, by Lemma 3.1, for each $\epsilon>0$ we have
$\log\#S_{\epsilon}=\limsup_{n\to\infty}\frac{1}{n}\log\#\hat{\mathcal{S}}_{n}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(\sigma,\xi).$
Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$ as $\epsilon\to 0$. Combining with the Theorem 1.3, one has
$\overline{\rm dim}_{B}(K)=\limsup_{\epsilon\to 0}\frac{\log\#S_{\epsilon}}{|\log\epsilon|}\leq\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm dim}_{B}(K).$
Finally, by the Corollary 1.5 we complete the proof.
Availability of data and materials: Not applicable.
Competing interests: The authors declare that they have no competing interests.
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University. The authors also thank Prof. Ercai Chen for many useful comments.

### v3

@@ 4.3 Generic points of full shift over compact metric space
Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
It is well-known that the system $(K^{\mathbb{Z}},\sigma)$ has the specification property. Moreover, it was proved in [VV17, Theorem 5] and [S22, Theorem 15] that,
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d),$ (4.4)
where $\overline{\rm dim}_{B}(K,d)$ is the upper box dimension of $K$ under the metric $d$.
Then, by Corollary 1.5, the metric mean dimensions of the generic points of the full shift over compact metric space are fully determined by the boxing dimension of the underlying space and mean Rényi information dimension of invariant measures.
Theorem 4.6.
Under the setting of Example 4.5, for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ one has
$\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d)$
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
Next, we present two examples of full shifts with the maximal metric mean dimension measure.
Example 4.7.
Consider the Example 4.4 again. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$, and define $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ as the product measure on $[0,1]^{\mathbb{Z}}$. Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\sigma,\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ25, Theorem 1.1], we have
@@
Endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$, and let $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ $1\leq i\leq l$ be contracting similitudes, that is,
$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}$
for some $r_{i}\in(0,1)$. One says that the iterated function system $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there is a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset,\ \text{for }i\neq j.$
Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure such that
  • (1)
    $K=\cup_{i=1}^{l}T_{i}(K)$;
  • (2)
    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, where $\alpha$ is the unique solution of equation
    $\sum_{i=1}^{l}r_{i}^{\alpha}=1,$
    and ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. $\alpha$ is called the similarity dimension of the IFS;
  • (3)
    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$;
  • (4)
    ${\rm supp}(\mu)=K$.
Many famous fractal sets, for instance, the middle thirds Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition.
Now take the attractor $K$ obtained above as in the Example 4.5. Then the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$ is invariant with respect to the left shift $\sigma$.
@@
Under the setting of Example 4.8, $\nu$ is the maximal metric mean dimension such that the three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha,$
where $S\in\{B,P,UC\}$
@@
Hutchinson [Hut81] showed that for any $z\in K$ and $\epsilon\in(0,1)$, one has
$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha},$ (4.5)
where $C>0$ is a constant independent of the choice of the point  $z\in K$ and  $\epsilon$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$ and $\epsilon\in(0,1)$, we define
[…]
Then the Bowen ball $B_{n}(x,\epsilon)$ is contained in $L_{n}(x,\epsilon)$ and $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5). Thus, one has
$\nu(B_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha},\ \forall x\in K^{\mathbb{Z}},$
which yields that $\overline{h}_{\nu}^{BK}(\sigma,\epsilon)\geq\alpha|\log\epsilon|-\log C.$ By the Shannon-McMillan-Breiman theorem, the following inequality is valid for any t.d.s. $(X,d,f)$:
$\overline{h}_{\mu}^{BK}(f,\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),\ \forall\mu\in M_{f}(X).$ (4.6)
Hence
[…]
which implies the desired result. ∎
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (No. 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University. The authors also thank Prof. Ercai Chen for many useful comments.